I did an experiment where I added 0.272 g of p-toluic acid (MW=136.1) and 0.357 g of N- bromosuccinimide (MW: 177.98), and 2 mg of benzoyl peroxide (MW: 242.2) and 2.5 mL chlorobenzene. The mass of the final product was 0.27 g. Benzoyl peroxide is an initiator so you do not need to determine the moles used in the experiment. Determine the limiting reagent, theoretical yield and the percentage yield.
Answer #1

Number of moles of p-toluic acid = 0.272g/136.1g/mol = 0.00199853 mol

Number of moles of N-bromosuccinimide = 0.357g/177.98g/mol = 0.0020058mol

The limiting reagent is p-toluic acid . (Answer)

Theoretical yield of the product (BrCH2C6H4COOH) = 0.00199853mol

Mass of the product formed (theoretical yield ) = 0.00199853mol× 215.044g/mol = 0.429772g = 0.43g (Answer)

Percentage yield = (0.27g/0.429772g)×100% = 62.824 %

= 62.8 %. (Answer)